Students, faculty, and staff line up for Tim Hortons. People are served at a rate of 3 per minute, on average.

Question 1: What is the probability that no one is served in the next minute? Round your answer to 4 decimal places.

Question 2 : What is the probability that 5 people are served in the next minute? Round your answer to 4 decimal places.

Question 3: What is the probability that fewer than 3 people are served in the next minute?
Answer #1
this is Poisson distribution with parameter λ=3

1)

probability that no one is served in the next minute =P(X=0)=e-3*30/0! =0.0498

2)

P(X=5)= {ex/x!}= 0.1008

3)

P(X<=2)= x=0x   {ex/x!}= 0.4232